Solve for y(t). dy/dt + 2x = et

dx/dt-2y= 1 +t

when x(0) = 1, y(0) = 2
Answer #1

1+2x=et dt - y =l+t Dy +2x = et w Du-2y = Citta 2D x + ²y = Det 2Dr - 4y= 214t) (0²741y=et_2-2t A.E of (3) is fm)=0 m²+4=0 m=5.3 x=tlet_by) ó et +2C Sin (et) – 2Ccos (2t). * ( 4ct +24870 (2+2–26.08(2t)) = Qet+ a sin(27)- Czcoscet) x (0)= 1 => 1 +2 +6